### v1

# Learning in Games with progressive hiding

Abstract
When learning to play an imperfect information game, it is often easier to first start with the basic mechanics of the game rules. For example, one can play several example rounds with private cards revealed to all players to better understand the basic actions and their effects. Building on this intuition, this paper introduces progressive hiding, an algorithm that learns to play imperfect information games by first learning the basic mechanics and then progressively adding information constraints over time. Progressive hiding is inspired by methods from stochastic multistage optimization such as scenario decomposition and progressive hedging. We prove that it enables the adaptation of counterfactual regret minimization to games where perfect recall is not satisfied. Numerical experiments illustrate that progressive hiding can achieve optimal payoff in a benchmark of emergent communication trading game.
1 Introduction
This paper shows how the learning process in games with imperfect information can be improved by relaxing the information constraints with penalty-like methods. For a game theorist, games are mathematical models to describe strategic interactions between agents (the players of the game). Games, in their standard acceptance, which usually supports this definition, are often referred to as the “Drosophila of Artificial Intelligence” [25]. This comparison comes from the fact that games provide Artificial Intelligence (AI) researchers with standardized, easily reproducible environments where they can test and benchmark various AI techniques. Games offer challenges that are complex enough to be interesting for AI research, but still have well-defined rules. One of the longstanding AI research questions about games is how to learn good strategies.
The field has advanced significantly due to breakthroughs where AI algorithms have outperformed human players in strategic games, illustrating the potent applications of AI in complex decision-making environments. Notable examples include Chess, Go, Poker and Stratego [39, 36, 26, 6, 7, 3, 30].
While Chess and Go are perfect information games, Stratego and Poker are imperfect information games, in the sense that the state of the game is not fully observed by the players. Hence players must account for the private information held by other players, as well as how their actions reveal their private information. Notably, this makes room for elements like bluffing [42]. Regret minimization at scale played a key role in recent achievements in AI for imperfect information games [26, 6, 7].
The AI community recently identified the card game Hanabi as a challenge for AI [1]. What differentiates Hanabi from the previously mentioned games is that it is a cooperative game 111the term same payoff games is also very often used: all players have the same objective. The difficulty comes from the uncertainty of the draws and from the fact that the players cannot communicate freely. The problem of pursuing a common objective with limited communication is not new, and spans several fields, in particular economics theory, control, game theory, and machine learning [44, 45, 22, 19, 27, 9, 16, 40, 15]. Witsenhausen showed, with his landmark counterexample [44], how imposing communication constraints on a problem drastically changes the nature of the problem.
We present a general framework for dealing with information in games. The framework is best described with the following real life observation: when teaching a child how to play a game — like Poker or Hanabi — with hidden information, one often starts with a few plays where all the cards are revealed, and then, as the child is learning, one starts hiding more and more information to make the game more challenging. This metaphor summarizes the idea of the main algorithm presented in this paper. Because of this metaphor— and also as a tribute to Rockafellar and Wets’ progressive hedging algorithm [35] [^2] See Setion 3 for related work.— we name the algorithm progressive hiding.
A common assumption in games is that players remember what they see and what they do in the precise order of occurrence. In his seminal work [18] on the tree representation of extensive form games, this assumption is what Kuhn referred to as perfect recall. One difficulty with a game such as Hanabi is that, if one wants to see the whole group of players as one big player, then this big player does not satisfy the perfect recall assumption because the one big player is not allowed to remember the private information of the other players. This absence of perfect recall prevents us from using key learning techniques [19, 21] such as counterfactual regret minimization (CFR) [46]. As we show, progressive hiding’s modification of the game allows recovering some notion of perfect recall, which makes the new problem amenable to CFR. In an effort to bridge ideas from decentralized control theory and computational game theory, we depart from the customary tree notations and introduce a hybrid between product games [15] and games on trees. This new, compact notation reveals quite practical for our purpose. While product games were fully formalized relatively recently, they are just the continuation of an idea that was hinted half a century ago by Witsenhausen [45], specifically as a way to deal with setting such as ours (communication constraints).
1.1 Contribution
Our main contribution – Theorem 2 – is to show how information relaxation can be made compatible with learning in games.
Section 2 introduces some background including the product game notation, that will prove practical for our purpose. Section 3 is an attempt to make a historical connection between algorithmic game theory and a stochastic programming technique called scenario decomposition. Section 4 presents a penalty method for games that can be analyzed under the lens of the proximal algorithms. This method is a first step toward our proposal, and brings the optimization tools that will be assembled with the learning ones in the second step. We then introduce in Section 5 progressive hiding (Algorithm 2), an algorithm that combines no-regret learning and information-relaxation. We show, in particular, how this approach allows us to recover the main property of CFR in an auxiliary game. We document numerical experiments performed on Trade Comm  [20, 40] in Section 6. Section 7 provides discussion on the limits of Algorithm 2. All proofs are provided in the appendix B.
2 Preliminary
This section is dedicated to background material on learning in games. There are different ways of modeling extensive form games [32], typically modeled as game trees. In this paper, we borrow ideas from the product form [15] as it better fits our needs for modeling the information structure. Games in product form represent information over a product set.
An important property of product games is that modifying the information structure of a product game can be done without modifying the product set over which this product game is defined. Such property is not satisfied by the tree representation (see Appendix E.2 on lumberjacking). Also, the product form provides a common framework for game theory and causality theory [14].
In what follows, if $S$ is a set and $s\in S$, we will use the notation $-s$ to refer to the set $\{s^{\prime}\in S,s^{\prime}\neq s\}$ when it is clear what $S$ is from the context. If $n$ is an integer, then we denote $[n]=\{1,2,\ldots,n\}$.
2.1 Extensive Game in Product Form
We first introduce a minimalistic version the product form 333without explicit references to the information algebras and the solution map for extensive form games, which is an alternative to the tree representation. The appendix contains more comprehensive information on the product game formalism. The reader can skim to Section 6 to see a simple instance of product game. In what follows, $P$ is the set of players in the game, $L$ is a positive integer that represents the maximum number of turns in the game (i.e., maximum length of any game ), and $W$ is a positive integer that represents the maximum number of legals actions at any time. We denote by $\Delta$ the $W$-dimensional simplex. We encode the available information to a player with an abstract, finite set $\mathbb{G}$. Otherwise said, $\mathbb{G}$ is a finite set that is used to encode the players’ information as the game proceeds. The randomness owned by the game itself (as opposed to the randomness controlled by the players), sometimes referred to as Nature’s moves, is encoded with a discrete probability space $(\Omega,\mathbb{P})$. The set $\Omega$ represents the space where random events take place. For example, in a card game, $\Omega$ would be the set of all possible orders of cards in the deck. In a dice game, it would be the set of all possible sequences of dice rolls during the game. There is a probability measure on $\Omega$, denoted by $\mathbb{P}$. For instance, in Poker, all possible orders of cards are equally probable. Last we define the product set as $\mathbb{H}=\Omega\times[W]^{L}$.
Given the primitives already introduced, a game in product form consists of a tuple $(\mathcal{P},\mathcal{A},\mathcal{X},r)$ where: $\mathcal{P}:[L]\to P$ is a map that indicates which player needs to play at stage $i\in[L]$; $\mathcal{A}:\mathbb{G}\to[W]$ is a map that encodes the number of available actions at each stage of the game; for all $i\in[L]$, a map $\mathcal{X}_{i}:\mathbb{H}\to\mathbb{G}$ that encodes the available information at each stage of the game, and that only depends on the first $i$ components of $\mathbb{H}$ (Nature’s move and the predecessors);444this assumption can be weakened, but helps the presentation the map $r:\mathbb{H}\to\mathbb{R}^{P}$ encodes the reward of the game, more precisely, $r_{p}$ is the reward of player $p$. For any element in $\mathbb{H}$, we denote by $h_{\emptyset}$ its first component (associated with Nature’s moves) and by $h_{i}$ the $(i+1)^{th}$ component, associated with the $i^{th}$ move of the players.
Compared to general extensive-form games [28, Chapter 11], Nature plays only once at the very beginning, and the random state is revealed progressively to the players. This choice of formalism is made without loss of generality and based on convenience for the problem considered in this paper.
2.2 Policies and Push-Forward Probability
Next, building on the primitives we just introduced, we can now describe how a product game is played by specifying the policies and the resulting push-forward probability. For a decision time $i\in[L]$, we denote by $\bar{\Lambda}_{i}$ the set of (admissible) policies $\mu_{i}$, where $\mu_{i}:\Omega\times[W]^{i-1}\to\Delta$ and $\operatorname{supp}{\mu_{i}(h)}\subset[\mathcal{A}\circ\mathcal{X}_{i}(h)]$, that is, the policy only selects admissible actions. Using a canonical extension, we can identify $\mu_{i}$ with an application from $\mathbb{H}\to\Delta$ that depends at most on Nature’s move and the first $i-1$ decisions. For $i\in[L]$, an implementable policy $\mu_{i}\in\bar{\Lambda}_{i}$ should also satisfy, for any histories $h$ and $h^{\prime}$ in the product space $\mathbb{H}$
$\mathcal{X}_{i}(h)=\mathcal{X}_{i}(h^{\prime})\implies\mu_{i}(h)=\mu_{i}(h^{\prime}).$ (1)
In words, an implementable policy depends solely on what player $\mathcal{P}(i)$ is supposed to know at the time of the decision. Relation (1) is the non-anticipativity constraint [37], and indicates that the policy $\mu_{i}$ should only depend on the information given by $\mathcal{X}_{i}$ to player $\mathcal{P}(i)$ at step $i$. If a policy is not implementable, it means that it needs some pieces of information that are not available at the time of the decision. We denote by $\Lambda_{i}$ the subset of policies from $\bar{\Lambda}_{i}$ that are implementable. Also, we denote by $\Lambda_{p}$ and $\bar{\Lambda}_{p}$ the sets of implementable and admissible policy profiles of player $p$ for $p\in[P]$. We similarly define $\Lambda$ and $\bar{\Lambda}$ to refer to policy profiles of all the players.
We refer to [45, 15] for more details on the following constructions. Given a policy profile $\mu$ and an element $\omega$ of $\Omega$, there is a unique element $h$ of $\mathbb{H}$ that satisfy $(\omega,\mu(h))=h$. Given a policy profile $(\mu_{i})_{i\in[L]}$, we can therefore construct the associated push-forward probability, denoted by $\mathbb{Q}_{\mu}$, on the product space $\mathbb{H}$. We denote by $\mathbb{E}_{\mu}$ the associated expectation operator. Last, we define a notation for modifying policies that allows for easy adjustment of specific components. Let $\mu$ be a base policy, and consider the operation of altering its $i$-th component to a new value $k$. This modification is denoted by $\mu{(i\to k)}$.  [^5] Formally, the modified policy $\mu(i\to k)$ is given by $\mu(i\to k)_{j}=\begin{cases}k&\text{if }j=i,\\ \mu_{j}&\text{otherwise}.\end{cases}$ This notation can be extended to accommodate multiple simultaneous modifications. For instance, changing the $i$-th component to $k$ and the $j$-th component to $l$ in $\mu$ is expressed as $\mu(i\to k)(j\to l)=\mu(i\to k,j\to l)$.
An $\epsilon$-Nash equilibrium is an admissible strategy profile $\mu$ so that no player can improve her outcome by more than $\epsilon$ from an unilateral deviation, that is, for any player $p$ and any other admissible strategy profile $\mu^{\prime}_{p}$, $\mathbb{E}_{\mu_{p},\mu_{-p}}[r_{p}(\textbf{h})]+\epsilon\geq\mathbb{E}_{\mu^{\prime}_{p},\mu_{-p}}[r_{p}(\textbf{h})].$ A Nash equilibrium is a 00-Nash equilibrium.
2.3 Perfect Recall and Information Maps
Perfect recall is a standard assumption in extensive-form games [38, 46]. An important property of perfect recall is that it implies the equivalence between mixed and behavioral strategies [18, 15]. In words, perfect recall means that the player has perfect memory of what they see and do and the precise order that information was revealed over the turns. In our context, perfect recall for player $p$ corresponds to the two conditions, for any $i<j$ such that $\mathcal{P}(i)=\mathcal{P}(j)=p$
$\bigg{(}\mathcal{X}_{i}(h)\neq\mathcal{X}_{i}(h^{\prime})\implies\mathcal{X}_{j}(h)\neq\mathcal{X}_{j}(h^{\prime})\bigg{)}\text{ and }\bigg{(}h_{i}\neq h_{i}^{\prime}\implies\mathcal{X}_{j}(h)\neq\mathcal{X}_{j}(h^{\prime})\bigg{)}.$ (2)
The first condition states that if the information known about $h$ and $h^{\prime}$ differs on turn $i$, then this must remain true at a future turn $j>i$. The second condition states that if the action taken at turn $i$ is different, then the player must remember this information at a later turn $j>i$. While the perfect recall assumption is realistic when the player corresponds to a single agent, this is less so when the player is used to model a team of several agents. The absence of perfect recall in a game prevents the use of several tools, for instance, backward induction, CFR [46, 19], and the equivalence of behavioral and mixed policies [18].
We say that an information map $\mathcal{X}_{i}^{+}$ is finer than an information map $\mathcal{X}_{i}$ if for any $h\in\mathbb{H}$ $\mathcal{X}_{i}(h)\neq\mathcal{X}_{i}(h^{\prime})\implies\mathcal{X}_{i}^{+}(h)\neq\mathcal{X}_{i}^{+}(h^{\prime})$. We say that $\Lambda^{+}\subset\bar{\Lambda}$ is induced by an information map profile $\mathcal{X}^{+}$ if it is the set of policy profiles from $\bar{\Lambda}$ that are implementable with respect to $\mathcal{X}^{+}$.
2.4 Projector
Let $\mu$ be any full support implementable policy profile, and $\tilde{\mathbb{H}}$ be the support of $\mathbb{Q}_{\mu}$. For $\mu_{0}\in\bar{\Lambda}$ of full support, we define $\mathrm{Proj}_{\mu_{0}}(\mu)$ as the element $\gamma$ of $\bar{\Lambda}$ such that
[…]
We have the following key properties (adapted from [35]):
Proposition 1.
$\forall\mu\in\bar{\Lambda},\mu\in\Lambda\iff\mathrm{Proj}_{\mu_{0}}(\mu)=\mu$.
Proposition 2.
$\mathrm{Proj}_{\mu_{0}}\circ\mathrm{Proj}_{\mu_{0}}=\mathrm{Proj}_{\mu_{0}}$.
Notably, proposition 1 provides an alternative representation of the non-anticipativity constraint (1).
2.5 No-regret learning in games
No-regret learning
The theory of online learning envision a decision maker that makes sequential decisions $x_{t}$ in a convex, compact set $\mathbb{X}$ at each epoch $t=1\ldots T$, afterwhat the (possibly adversarial) environment outputs a concave reward $\ell_{t}$. An important metric of success for an algorithm is the regret, defined as $\max_{x\in\mathbb{X}}\sum_{t=1}^{T}\ell_{t}(x)-\sum_{t=1}^{T}\ell_{t}(x_{t}).$ A regret minimization algorithm is one where this regret grows at a rate of $o(T)$. An algorithm with this property is called Hannan-consistent. Regret minimization algorithms play a central role in optimization and game theory, because they can be used to maximize a function or find approximate Nash equilibrium [2]. For example, in the zero-sum two-player setting, one can find $\epsilon$-Nash equilibrium by repeating the game with two Hannan-consistent learners.
CFR
[…]
3 Related Works
In this section, we attempt to draw a connection between algorithmic game theory and a stochastic programming technique called scenario decomposition. As argued in [35] a common methodology for addressing uncertainty involves initially adopting a deterministic simplification. Practitioners first solve the problem for some known scenarios, thereby establishing a baseline of deterministic solutions. Subsequently, these solutions are aggregated to incorporate uncertainty. This approach simplifies complex problems by breaking them down into more manageable components.
In game theory, this type of approach is sometimes branded as Perfect Information Monte Carlo search (PIMC search) [24]). Motivated by the unexpected success of the methods in Bridge [12], Skat [8] and Hearts [43], the authors of [24], discuss conditions for PIMC to work. While the theoretical limit of the PIMC approach has been recognized very early [10], it still achieves state of the art performance in games such Skat and Bridge. Further improvements of PIMC are proposed in [11, 41].
Paraphrasing the authors of [35], the aim of progressive hedging is to provide "a rigorous algorithmic procedure for determining (…) a policy in response to any weighting of the scenarios". The algorithm iteratively tweaks the rewards in each scenarios so that at some point, the procedure finds a policy that is adapted to the uncertainty filtration. The algorithm leverages the ideas of augmented Lagrangian and proximal algorithms, and enjoys theoretical guaranties under some convexity assumptions. Progressive hedging was recently adapted to variational inequalities [34]. In the same vein, the recent monograph [37] presents Lagrangian methods applied to non-anticipativity constraints in the context of stochastic programming. The discretization of the random space and the introduction of stochastic dual variables allows decomposing stochastic programs per scenarios. The dual interpretation comes with important results from convex analysis [33].
Crucially, in with these stochastic programming techniques, the decomposition is done per scenario, not per agent. To apply online learning techniques, we would prefer to obtain a decomposition per agent. Furthermore, in stochastic optimization, a sole decision maker optimizes jointly all the decisions. By contrast, in games with imperfect information, players do not observe the same information, and the action of a player can influence what the other players observe.
4 Information Relaxation
In this section, we combine the penalty method and proximal algorithms [29] to produce a player best reply, that is, a policy that is optimal given the policy of the other players. While the resulting Algorithm 1 might be of independent interest, it serves as a guiding principle for the design of progressive hiding, presented in the next section. We first reduce to the case where $|P|=1$ by fixing the policies of all players but one. Let $\tilde{\Lambda}\subset\bar{\Lambda}$ induced by an information map profile $\mathcal{\tilde{X}}$ such that $\Lambda\subset\tilde{\Lambda}$. Instead of restricting our search to implementable policies, we could optimize over $\tilde{\Lambda}$. However, the optimal policy would very likely not satisfy the non-anticipativity constraint. Using Proposition 1, we suggest to favorize implementability through a penalty term $-\lambda||\mu-\mathrm{Proj}(\mu)||^{2}$, where $\mathrm{Proj}$ is a projector as introduced in Section 2.4 induced by a policy $\mu_{0}$ and |||||$\cdot$||| | ⋅ | | is the reweighted $L_{2}$ norm associated with $\mu_{0}$. This idea drives us to a relaxed version of the problem of finding an implementable best reply. Indeed, setting $\mathcal{L}(\mu,\lambda)=\mathbb{E}_{\mu}[r_{p}(\textbf{h})]-\lambda||\mu-\mathrm{Proj}(\mu)||^{2},$ we propose to solve the relaxed problem
$\max_{\mu\in\tilde{\Lambda}}\mathcal{L}(\mu,\lambda).$ (4)
Observe that, because of the penalty, the criterion is unusual and might be impractical for many game solving approaches. A first step to address this difficulty is Theorem 1, which states that one can produce a local optimum by composing two proximal steps, which can be analyzed under the lens of the Majorize-Minimization Algorithm (Algorithm 1). Progressive hiding builds on this idea by adding a linearization step as in [47] for online convex problems.
Input: $\lambda>0$, $\mu^{(0)}\in\bar{\Lambda}$, $T\in\mathbb{N}$
Initialization: $\mu=\mu^{(0)}$
for $t\leftarrow 1$ to $T$ do
       $\gamma\leftarrow\text{Proj}(\mu)$
        // projection step
       $\mu\leftarrow\arg\max_{\mu\in\tilde{\Lambda}}\mathbb{E}_{\mu}[r_{p}(\textbf{h})]-\lambda||\mu-\gamma||^{2}$
        // proximal step
endfor
return $(\mu,\gamma)$
Algorithm 1 Resolution by Information Relaxation
Theorem 1.
Fix $\lambda>0$, $\mu^{0}\in\bar{\Lambda}$ and denote by $(\mu^{T},\gamma^{T})$ the output of Algorithm 1, then $\mathcal{L}(\mu^{T},\lambda)$ is non-decreasing and admits a limit.
Proposition 3.
We have the relation $\max_{\mu\in\Lambda}\mathbb{E}_{\mu}[r_{p}(\textbf{h})]\leq\max_{\mu\in\bar{\Lambda}}\mathbb{E}_{\mu}[r_{p}(\textbf{h})]-\lambda||\mu-\mathrm{Proj}(\mu)||^{2}$.
The type of upper bound of Proposition 3 is often used to estimate duality gap in operations research [4]. Observe that when $\lambda$ is small, Problem (4) becomes closer to the problem of finding an optimal policy in $\tilde{\Lambda}$, so the couple $(\mu,\gamma)$ produced by Algorithm 1 might not satisfy $\mu\approx\gamma$ and $\mu$ might be far from implementable. By contrast, if $\lambda$ is large, one expects $\mu$ and $\gamma$ to be close, however, because of the proximal term, in each iteration only solutions very close to the previous steps will be considered, and we might get stuck with a local optima. In practice, many strategies could be used to solve the $\arg\max$ step in Algorithm 1. In particular, if the auxiliary game (with $\tilde{\Lambda}$ instead of $\Lambda$) satisfies perfect recall, then one can rely on backward induction. The next section, which presents our main contribution, is guided by the idea of combining the philosophy of Algorithm 1 with regret minimization approaches.
5 Progressive Hiding
We next present an algorithm — progressive hiding — that replicates the philosophy of information relaxation in the framework of no-regret learning by introducing an auxiliary game. This section also contains the article principal contribution, which is Theorems 2. We suppose the players in $-p=P\setminus\{p\}$ play according to a policy profile $\mu^{t}_{-p}$ at stage $t$.
5.1 Algorithm
We suppose we have access to a class of low-regret algorithms that implements two functions: Observe (to get the realization of vector of linear reward) and Decide (to output a probability distribution on the decision). Progressive hiding requires at each time $t$ a non-negative penalty parameter $\lambda^{t}$ for the non-anticipativity constraint violation. This sequence of parameters $\lambda^{t}$ might be time adaptive to allow for more flexibility at the beginning, but still shift toward implementability at the end of the learning process. Progressive hiding also requires an information map $\tilde{\mathcal{X}}$, which is a refinement (cf. Section 2.3) over the information map $\mathcal{X}$. For example, $\tilde{\mathcal{X}}$ can inform the player about the hand of its opponents, or about what other teammates saw in the previous turns. The finer the information map $\tilde{\mathcal{X}}$, the more additional information is provided in the auxiliary games. Notably, ensuring that perfect recall is satisfied in the auxiliary game seems to be a fruitful direction, as indicated by Theorem 2. Last, progressive hiding also requires a sequence of projectors $\text{Proj}^{t}$ as defined in Section 2.4. We set $\mathrm{Proj}^{t}=\mathrm{Proj}_{\mu^{t}_{p},\mu^{t}_{-p}}$. We denote by $I_{p}$ the elements $i$ of $[L]$ such that $\mathcal{P}(i)=p$, and we set $G_{i}=\tilde{\mathcal{X}}_{i}(\mathbb{H})$, for $i\in I_{p}$. We introduce the notation $\mathbb{E}_{\mu}[\textbf{r}_{p}|g]=\mathbb{E}_{\mu}[r_{p}(\textbf{h})|g\in\{\tilde{\mathcal{X}}_{i}(\textbf{h}),i\in I_{p}\}]$. Last, we introduce the time-dependent, random penalty, for $i\in I_{p}$ and $h\in\tilde{\mathbb{H}}$
$\ell^{t}_{i}(\mu_{i}(h),\tilde{\mathcal{X}}_{i}(h))=\lambda^{t}||\mu_{i}(\tilde{\mathcal{X}}_{i}(h))-\text{Proj}^{t}(\mu^{t}_{p})_{i}(\tilde{\mathcal{X}}_{i}(h))||^{2}.$ (5)
The right-hand side of Equation (5) makes sense because (1) $\mu_{i}$ is a function of $\tilde{\mathcal{X}}_{i}(h)$, (2) $\text{Proj}^{t}(\mu^{t}_{p})$ is a function of $\mathcal{X}_{i}(h)$ and $\tilde{\mathcal{X}}$ is finer than $\mathcal{X}$. Progressive hiding is displayed in Algorithm 2.
Initialization: Initialize $\text{RegrMin}[i,g]$ for $(i,g)\in I_{p}\times G_{i}$
for $t\leftarrow 1$ to $T$ do
       foreach $(i,g)\in I_{p}\times G_{i}$ do
             $\mu^{t}_{i}(g)\leftarrow\text{RegrMin}[i,g].\textsc{Decide}()$
              // local no-regret decisions
       end foreach
      $\gamma\leftarrow\textsc{Proj}^{t}(\mu^{t}_{p})$
        // Projection step
       foreach $(i,g)\in I_{p}\times G_{i}$ do
             $\theta_{i,g}\leftarrow(\varrho_{t}[i,g][\delta_{d}])_{d\in[\mathcal{A}_{i}(g)]}$
              // reward in the auxiliary game
             $\text{RegrMin}[i,g].\textsc{Observe}\left(\theta_{i,g}\right)$
              // feed the local learners
       end foreach
endfor
return $\gamma$
Algorithm 2 Progressive Hiding
Interpretation
Suppose that at stage $t$ of the algorithm, the policy is $\mu^{t}_{i}$ for $i\in I_{p}$. Then for $i\in I_{p}$ and $g\in G_{i}$, the local reward vector that is fed to the regret minimizer is $\mathbb{E}_{\delta_{k},\mu_{-i}^{t},\mu_{-p}^{t}}[r_{p}(\textbf{h})|\tilde{\mathcal{X}}_{i}(\textbf{h})=g]$ with $k$ in $[\mathcal{A}(g)]$. We add a term proportional to the gradient of $-||\gamma_{i}(g)-\mu^{t}_{i}(g)||^{2}$ so that we optimize for the penalized criterion (4). It should be noted that thanks to the projection step, the algorithm output an implementable policy no matter the choice of $\lambda$ and $\tilde{\mathcal{X}}$. While the rational is different, we still note that the modification of a game payoff with a proximal term that makes the full criterion policy-dependent is also observed in [31]. In our case, however, the proximal term contains a projection, and is used to push the solution toward implementability.
5.2 Properties
Let $\mu^{t}$ be the iterates of progressive hiding defined in Algorithm 2. We denote respectively by
$\varrho_{t}(\mu_{p})=\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}r_{p}(\textbf{h})-\sum_{i\in I_{p}}\ell^{t}_{i}(\mu_{i}(\textbf{h}),\tilde{\mathcal{X}}_{i}(\textbf{h}))\bigg{]},\quad\mbox{and}\quad R^{T}=\nicefrac{{1}}{{T}}\max_{\mu\in\tilde{\Lambda}}\sum_{t=1}^{T}\varrho_{t}(\mu_{p})-\varrho_{t}(\mu^{t}_{p}),$
the new criteria for the player of interest at a given time step $t$, and the average regret $R^{T}$ in the auxiliary game with relaxed information constraints and penalty for violation. Also, for $i\in I_{p}$ and $g\in G_{i}$, we set
$\varrho_{t}[i,g][\mu_{i}]=\mathbb{E}_{\mu^{t}_{p}(i\to\mu_{i}),\mu_{-p}}\Bigg{[}r_{p}(\textbf{h})-\lambda^{t}\bigg{(}2\langle\mu_{i}^{t}(\textbf{h})-\textsc{Proj}^{t}(\mu^{t}_{p})_{i}(\textbf{h})\mid\mu_{i}\rangle\bigg{)}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(\textbf{h}),\tilde{\mathcal{X}}_{i^{\prime}}(\textbf{h}))\bigg{|}g\Bigg{]},$
which corresponds to the local criterion optimized by progressive hiding, and $R^{T}_{loc}(i,g)=\nicefrac{{1}}{{T}}\max_{\mu_{i}\in\tilde{\Lambda}_{i}}\sum_{t=1}^{T}\varrho_{t}[i,g](\mu_{i})-\varrho_{t}[i,g](\mu^{t}_{i})$ the local regret associated with this local criterion. We also set $R^{T,+}_{loc}(i,g)=\max\left(R^{T}_{loc}(i,g),0\right)$.
Theorem 2.
Suppose $\tilde{\mathcal{X}}$ satisfies the perfect recall condition 2. Then the overall regret is bounded according to the relation $R^{T}\leq\sum_{i\in I_{p}}\sum_{g\in G_{i}}R^{T,+}_{loc}(i,g)$.
Otherwise said, Theorem 2 states that CFR can be applied in the auxiliary game. However, note that the regret $R^{T}$ is defined with respect to the auxiliary game payoff. Mirroring Theorem 1 for Algorithm 1, Theorem 2, with this bound on the regret in the auxiliary game, quantifies what Algorithm 2 seeks.
The next property (which does not require perfect recall to hold) allows sizing the distance between $\mu^{t}$ and the implementable set.
Proposition 4.
$\nicefrac{{1}}{{T}}\sum_{t=1}^{T}\left(\lambda^{t}\mathbb{E}_{\mu_{p}^{t},\mu^{t}_{-p}}\sum_{i\in I_{p}}\bigg{[}||\mu^{t}_{i}(\textbf{h})-\text{Proj}^{t}(\mu^{t}_{p})_{i}(\textbf{h})||^{2}\bigg{]}\right)\leq R^{T}+2.||r||_{\infty}.$
While we leave the choice of $\lambda^{t}$ open, we report from our preliminary experiments that selecting a constant value with a grid search seems a viable strategy.
An important assumption to guarantee low-regret with counterfactual regret minimization is that the player (or the group of player) satisfies perfect recall. In practice, to cope with games of large size, the algorithm is often applied with an approximation of the game representation, which implies that perfect recall might not hold. Similarly, here, while we can provide guaranty when the relaxation induces perfect recall to hold, we also believe that for applications, an approximation of the perfect recall could also be envisioned for scalability concerns.
6 Experiments
This section reports on preliminary numerical experiments that complement the theoretical findings. We run two versions of progressive hiding on Trade Comm [20] because the game is elementary to solve for a human, but appears to be quite challenging for learning algorithms [40]. Furthermore, the game is parametrized by two parameters, which allows us to test different settings. The code is provided in the Supplementary Material. Of independent interest, the code closely aligns with the product game description, hence despite their equivalence in our context, the tree and product representations lead to distinct coding paradigms.
Figure 1: Learning outcomes ($\mathbb{E}_{\gamma^{t}}[r_{t}(\textbf{h})]$) distribution for the three information map baseline, recall and cheated on Trade Comm with parameter $(m,n)=(2,2)$ (left) and $(m,n)=(3,2)$ (right).
Trade Comm is a common-payoff communication game. Each of the two players randomly receives one of several items in $[n]$ ($s_{1}$ and $s_{2}$). The first player communicates by choosing one out of several possible messages in $[m]$, denoted by $m_{1}$, which the second player observes. Then, the second player also selects a message (in $[m]$), denoted by $m_{2}$, for the first player to see. Both players then secretly request a trade involving one of the possible item combinations ($(d_{1}^{1},d_{2}^{1})$ and $(d_{1}^{2},d_{2}^{2})$). A trade is successful if both players request to exchange their items for each other’s. They both earn a point if the trade works out, and none if it doesn’t. Despite seeming straightforward, Trade Comm effectively highlights the challenges in same payoff games. We showcase that information relaxation offers a range of strategies through the choice of the information map $\tilde{\mathcal{X}}$ and the penalty schedule $\lambda^{t}$.
Formally, $P=1$, $(s_{1},s_{2})\in\Omega=[n]\times[n]$, $\mathbb{P}$ is the uniform distribution on $\Omega$, the product space is
$\mathbb{H}=\underbrace{([n]\times[n])}_{\Omega}\times\underbrace{[m]}_{\text{first message}}\times\underbrace{[m]}_{\text{second message}}\times\underbrace{([n]\times[n])}_{\text{first trade request }}\times\underbrace{([n]\times[n])}_{\text{second trade request }}.$
The information map $\mathcal{X}$ is $\mathcal{X}_{1}(h)=s_{1}$, $\mathcal{X}_{2}(h)=s_{2}$, $\mathcal{X}_{3}(h)=(s_{1},m_{1},m_{2})$, $\mathcal{X}_{4}(h)=(s_{2},m_{2},m_{1})$, the available actions map $\mathcal{A}$ is $\mathcal{A}_{1}(h)=[m]$, $\mathcal{A}_{2}(h)=[m]$, $\mathcal{A}_{3}(h)=[n]\times[n]$, $\mathcal{A}_{4}(h)=[n]\times[n]$, and, last, the reward is $r_{p}(h)=[d^{1}_{1}=s_{1}]\cdot[d^{2}_{2}=s_{1}]\cdot[d^{2}_{1}=s_{2}]\cdot[d^{1}_{2}=s_{2}]$. We can check that perfect recall is not satisfied for $n>1$.
We envision three possible information maps: the "original" information map $\mathcal{X}$, the "cheater" information map that reveals the private information, but does not record the message sent. This type of imperfect recall can be desired for computational efficiency in the same spirit as game abstractions are used: $\mathcal{X}_{1}^{cheat}(h)=(s_{1},s_{2})$, $\mathcal{X}_{2}^{cheat}(h)=(s_{2},s_{1})$, $\mathcal{X}_{3}^{cheat}(h)=(s_{1},s_{2},m_{2})$, $\mathcal{X}_{4}^{cheat}(h)=(s_{2},s_{1},m_{1})$, and a perfect recall information map, that, as it names suggest, ensures perfect recall: $\mathcal{X}_{1}^{PR}(h)=(s_{1})$, $\mathcal{X}_{2}^{PR}(h)=(s_{1},s_{2},m_{1})$, $\mathcal{X}_{3}^{PR}(h)=(s_{1},s_{2},m_{1},m_{2})$, $\mathcal{X}_{4}^{PR}(h)=(s_{1},s_{2},m_{1},m_{2},d_{1}^{1},d_{2}^{1})$.
Appendix F contains more details on the experiments. We compare the three information maps for $(n,m)=(2,2)$ and $(n,m)=(3,2)$ in Figure 1. The dashed lines correspond to the 10% quantiles, and the full line to the average value, over 100 runs with randomized initial policy. We see that the methods relying on progressive hiding beat the baseline on the two examples. Progressive hiding reaches the optimal payoff more often and faster than the baseline. It is notable that even cheated, which consists in an information relaxation not big enough to satisfy perfect recall, still induces a clear improvement over the baseline.
7 Discussion
In this article, we show how information relaxation can be made compatible with learning in games. We build on no-regret learners to propose progressive hiding. We show that the algorithm is principled and showcase promising experimental results.
Limitation
The main limitation of progressive hiding is its local nature, which stems from our definition of regret. Another limitation of progressive hiding is its scalability, but we can be optimistic in the possibility to combine the method with Monte Carlo sampling [21] and deep learning approaches [5, 13]. Last, a question that is not addressed in this work is how to choose $\lambda^{t}$.
Further work
A question that is not fully answered in this work is the choice of information map to perform the information relaxation. It seems intuitive that this question relates to the literature on game abstractions. Also, while we introduced information maps (i.e., functions) to encode the player’s knowledge for the sake of simplicity, it seems that information fields [9] because they have a lattice and algebraic structure, would be a better mathematical tool if one wants to go further. Typically, an information field of interest is the smallest field greater than the original field that ensures that perfect recall is satisfied. In this direction, it would be interesting to know if there exist informational properties other than perfect recall that allow for decomposition results.
Appendix
Appendix A Preliminary results: extension of CFR
A.1 Introduction
In order to prove our main claims, we first prove an extension of CFR [46] where a term is added to the game theoretical objective. The result and the derivations are of independent interest. Note that this is done with the formalism of product games. In what follows, we suppose perfect recall is satisfied for a player $p\in P$ of interest (for the information map $\tilde{\mathcal{X}}_{p}$). Without loss of generality because we are under perfect recall, we suppose the intersection of the $G_{i}$ to be empty. If it exists, we denote by $i+_{p}1$ the next element after $i$ in $I_{p}$ for $i\in I_{p}$. We denote by $>_{p}i$ all the elements that come after $i$ in $I_{p}$, and by $\geq_{p}i$ the union of $>i$ and the singleton $\{i\}$. We then write $j>_{p}i$ if $j\in>_{p}i$. We define the notations $<_{p}$ and $\leq_{p}$ similarly. We set $\text{Succ}[i,g]=\cup_{i^{\prime}>_{p}i}\tilde{\mathcal{X}}_{i^{\prime}}(\tilde{\mathcal{X}}_{i}^{-1}(g))$ for $i\in I_{p}$ and $g\in G_{i}$. Let $r_{p}$ be a payoff defined on $\mathbb{H}$ and $\ell^{t}_{i}$ some penalties from $\Delta\times\mathbb{G}$ to $\mathbb{R}$. Let $\mu^{t}_{p}\in\Lambda_{p}$ and $\mu^{t}_{-p}\in\Lambda_{-p}$ for all $t$. We make the very practical observation that for $i\in I_{p}$ and $g\in G_{i}$, the quantity $\mathbb{E}_{\mu_{p},\mu_{-p}}[\textbf{r}_{p}|g]$ does not depend on $\mu_{j}$ for $j<_{p}i$. This allows us to use the notation $\mathbb{E}_{\mu_{>_{p}i},\mu_{-p}}[\textbf{r}_{p}|g]$.
A.2 Problem statement
At time $t$, let be the criteria
$\mathbb{E}_{\mu_{p}\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i\in I_{p}}\ell^{t}_{i}(\mu_{i}(h),\tilde{\mathcal{X}}_{i}(h))\bigg{]}.$
We are interested in understanding the relations between the following two quantities:
$R^{T}$ $=\max_{\mu_{p}\in\Lambda_{p}}\nicefrac{{1}}{{T}}\sum_{t=1}^{T}\Bigg{(}\mathbb{E}_{\mu_{p}\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i\in I_{p}}\ell^{t}_{i}(\mu_{i}(h),\tilde{\mathcal{X}}_{i}(h))\bigg{]}-\mathbb{E}_{\mu_{p}^{t}\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i\in I_{p}}\ell^{t}_{i}(\mu_{i}^{t}(h),\tilde{\mathcal{X}}_{i}(h))\bigg{]}\Bigg{)},$
$R_{imm}^{T}[i,g]$ $=\nicefrac{{1}}{{T}}\max_{\mu_{i}\in\Lambda_{i}}\sum_{t=1}^{T}\bigg{(}\ell^{t}_{i}(\mu_{i}^{t}(g),g)-\ell^{t}_{i}(\mu_{i}(g),g)+$ (6)
$\mathbb{E}_{\mu_{i},\mu^{t}_{-i},\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]},$
for $i\in I_{p}$ and $g\in G_{i}$. More precisely, our goal is to show the following relation $R^{T}\leq\sum_{i\in I_{p}}\sum_{G_{i}}R_{imm}^{T,+}[i,g]$, (with the shorthand $R_{imm}^{T,+}[i,g]=\max(0,R_{imm}^{T}[i,g])$) which looks like the relation on which CFR is built.
To do so, we first introduce the following intermediate quantity:
$R_{full}^{T}[i,g]$ $=\max_{\mu_{p}\in\Lambda_{\geq_{p}i}}\nicefrac{{1}}{{T}}\sum_{t=1}^{T}\left(\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\right).$
Observe that aside from the existence of an ordering, we make no specific assumptions on the information available to the other players at decision time. Remember that $\tilde{\mathcal{X}}_{p}$ is a refinement over $\mathcal{X}_{p}$.
A.3 Derivation
We start with the following lemma, akin to backward induction.
Lemma 1.
For $i\in I_{p}$, $g\in G_{i}$,
[…]
$+\mathbb{E}_{\mu_{i},\mu^{t}_{-i},\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}$
$+\max_{\mu^{+}\in\Lambda_{>_{p}i}}\sum_{t=1}^{T}\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{i},\mu^{t}_{-i},\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\bigg{)}.$
Proof.
By definition of $R_{full}^{T}[i,g]$,
$R_{full}^{T}[i,g]=\max_{\mu_{p}\in\Lambda_{\geq_{p}i}}\nicefrac{{1}}{{T}}\sum_{t=1}^{T}\left(\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\right)$
by property of the $\max$ ($\max_{x,y}=\max_{x}\max_{y}$),
$=\max_{\mu_{i}\in\Lambda_{i}}\max_{\mu^{+}\in\Lambda_{>_{p}i}}\nicefrac{{1}}{{T}}\sum_{t=1}^{T}\left(\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\right)$
grouping toguether what does not depend on $\mu^{+}$,
$=\nicefrac{{1}}{{T}}\max_{\mu_{i}\in\Lambda_{i}}\sum_{t=1}^{T}\left(-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}+\max_{\mu^{+}\in\Lambda_{>_{p}i}}\sum_{t=1}^{T}\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}\geq_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\right)$
by the fact that $\tilde{\mathcal{X}}_{i}(h)=g$, and using the abuse of notation $\mu_{i}(h)=\mu_{i}(g)$,
$=\nicefrac{{1}}{{T}}\max_{\mu_{i}\in\Lambda_{i}}\sum_{t=1}^{T}\bigg{(}\ell^{t}_{i}(\mu_{i}^{t}(g),g)-\ell^{t}_{i}(\mu_{i}(g),g)-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}$
$+\max_{\mu^{+}\in\Lambda_{>_{p}i}}\sum_{t=1}^{T}\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\bigg{)}$
adding and removing the term $\mathbb{E}_{\mu_{i},\mu^{t}_{-i},\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}$
$=\nicefrac{{1}}{{T}}\max_{\mu_{i}\in\Lambda_{i}}\sum_{t=1}^{T}\bigg{(}\ell^{t}_{i}(\mu_{i}^{t}(g),g)-\ell^{t}_{i}(\mu_{i}(g),g)$
$+\mathbb{E}_{\mu_{i},\mu^{t}_{-i},\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{p}^{t},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}$
$+\max_{\mu^{+}\in\Lambda_{>_{p}i}}\sum_{t=1}^{T}\mathbb{E}_{\mu_{p},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{i},\mu^{t}_{-i},\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\bigg{)},$
which is the relation claimed in the lemma. ∎
For the next result, we use the shorthands $R_{imm}^{T,+}[i,g]=\max(0,R_{imm}^{T}[i,g])$ and $R_{full}^{T,+}[i,g]=\max(0,R_{full}^{T}[i,g])$ We use Lemma 1 to obtain the following bound on $R_{full}^{T}[i,g]$.
Lemma 2.
For $i\in I_{p}$, $g\in G_{i}$,
$R_{full}^{T,+}[i,g]\leq R_{imm}^{T,+}[i,g]+\sum_{g^{\prime}\in succ[i,g]}R_{full}^{T,+}[i+_{p}1,g^{\prime}])$
Proof.
Starting from Lemma 1 we use the fact that $\max_{x}f(x)+g(x)\leq\max_{y}f(y)+\max_{z}g(z)$ and the definition of $R_{imm}^{T}[i,g]$ to get
$R_{full}^{T}[i,g]\leq R_{imm}^{T}[i,g]$
$+\nicefrac{{1}}{{T}}\max_{\mu_{i}\in\Lambda_{i}}\max_{\mu^{+}\in\Lambda_{>_{p}i}}\sum_{t=1}^{T}\mathbb{E}_{\mu_{i},\mu^{+},\mu_{-p}^{t}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}-\mathbb{E}_{\mu_{i},\mu^{t}_{-i},\mu^{t}_{-p}}\bigg{[}\textbf{r}_{p}-\sum_{i^{\prime}>_{p}i}\ell^{t}_{i^{\prime}}(\mu_{i^{\prime}}^{t}(h),\tilde{\mathcal{X}}_{i^{\prime}}(h))|g\bigg{]}\bigg{)}.$
we bound the RHS by applying $\max(0,x)$ on the term and using the fact that perfect recall holds for player $p$
$\leq R_{imm}^{T,+}[i,g])+\sum_{g^{\prime}\in succ[i,g]}R_{full}^{T,+}[i+_{p}1,g^{\prime}],$
From this relation, we deduce $R_{full}^{T,+}[i,g]\leq R_{imm}^{T,+}[i,g]+\sum_{g^{\prime}\in succ[i,g]}R_{full}^{T,+}[i+_{p}1,g^{\prime}])$, QED.
Now we can state our modified CFR lemma.
Lemma 3.
$R^{T}\leq\sum_{i\in I_{p}}\sum_{G_{i}}R_{imm}^{T,+}[i,g]$
Proof.
we combine an induction, Lemma 2 and the fact that by assumption, perfect recall holds (which implies that the successors’ appearances are unique). ∎
Appendix B Proofs of Section 2
Proposition (1).
$\forall\mu\in\bar{\Lambda},\mu\in\Lambda\iff\mathrm{Proj}_{\mu_{0}}(\mu)=\mu$
Proof.
  • ()($\implies$)( ⟹ )
    If $\mu\in\Lambda$, then by definition of $\Lambda$, it should satisfy relation (1). Hence in equation (3), we get $\gamma_{i}(h)=\mathbb{E}_{\mu_{0}}[\mu_{i}(\textbf{h'})|\mathcal{X}_{i}(\textbf{h'})=\mathcal{X}_{i}(h)]=\mathbb{E}_{\mu_{0}}[\mu_{i}(h)|\mathcal{X}_{i}(\textbf{h'})=\mathcal{X}_{i}(h)]=\mu_{i}(h)$. Hence $\mathrm{Proj}_{\mu_{0}}(\mu)=\mu$.
[…]
    For any $i\in[L]$ and $h$ and $\tilde{h}$ in $\mathbb{H}$ such that $\mathcal{X}_{i}(h)=\mathcal{X}_{i}(\tilde{h})$, then $\gamma_{i}(h)=\mathbb{E}_{\mu_{0}}[\mu_{i}(\textbf{h'})|\mathcal{X}_{i}(\textbf{h'})=\mathcal{X}_{i}(h)]=\mathbb{E}_{\mu_{0}}[\mu_{i}(\textbf{h'})|\mathcal{X}_{i}(\textbf{h'})=\mathcal{X}_{i}(\tilde{h})]=\gamma_{i}(\tilde{h})$. Hence, $\mathrm{Proj}_{\mu_{0}}(\mu)$ satisies relation (1), and thus belong to $\Lambda$. Finally, since $\mathrm{Proj}_{\mu_{0}}(\mu)=\mu$, we conclude that $\mu\in\Lambda$.
Proposition (2).
$\mathrm{Proj}_{\mu_{0}}\circ\mathrm{Proj}_{\mu_{0}}=\mathrm{Proj}_{\mu_{0}}$
Proof.
Take $\mu\in\bar{\Lambda}$, from the second step of the proof of Proposition 1, we see that $x:=\mathrm{Proj}_{\mu_{0}}(\mu)\in\Lambda$, therefore, by 1, $\mathrm{Proj}_{\mu_{0}}(x)=x$
Appendix C Proof of Section 4
Theorem (1).
Fix $\lambda>0$, $\mu^{0}\in\bar{\Lambda}$ and denote by $(\mu^{T},\gamma^{T})$ the output of Algorithm 1, then $\mathcal{L}(\mu^{T},\lambda)$ is non-decreasing and admits a limit.
Proof.
Set $f(\mu)=\mathcal{L}(\mu,\lambda)=\mathbb{E}_{\mu}[r_{p}(\textbf{h})]-\lambda||\mu-\text{Proj}(\mu)||^{2}$ and $g(\mu,\gamma)=\mathbb{E}_{\mu}[r_{p}(\textbf{h})]-\lambda||\mu-\gamma||^{2}$, then
$f(\mu^{T-1})=g(\mu^{T-1},\gamma^{T-2})\leq g(\mu^{T-1},\gamma^{T-1})\leq g(\mu^{T},\gamma^{T-1})=f(\mu^{T}).$
Therefore, $T\to f(\mu^{T})$ is non-decreasing and bounded, hence it converges. ∎
Proposition (3).
$\max_{\mu\in\Lambda}\mathbb{E}_{\mu}[r_{p}(\textbf{h})]\leq\max_{\mu\in\bar{\Lambda}}\mathbb{E}_{\mu}[r_{p}(\textbf{h})]-\lambda||\mu-\mathrm{Proj}(\mu)||^{2}$
Proof.
This is a direct consequence from the fact that $\Lambda\subseteq\bar{\Lambda}$ and $\mathrm{Proj}(\mu)=\mu$ on $\Lambda$. ∎
Appendix D Proof of section 5
Theorem (2).
Suppose $\tilde{\mathcal{X}}$ satisfies the perfect recall condition 2. Then the overall regret is bounded according to the relation $R^{T}\leq\sum_{i\in I_{p}}\sum_{g\in G_{i}}R^{T,+}_{loc}(i,g)$.
Proof.
By Lemma 3, $R^{T}\leq\sum_{i\in I_{p}}\sum_{g\in G_{i}}R^{T,+}_{imm}(i,g)$ (where $R_{imm}^{T}$ has been defined in relation (6)). We then observe that by convexity of the squared norm, $R^{T,+}_{imm}(i,g)\leq R^{T,+}_{loc}(i,g)$. ∎
Proposition (4).
$\nicefrac{{1}}{{T}}\sum_{t=1}^{T}\left(\lambda^{t}\mathbb{E}_{\mu_{p}^{t},\mu^{t}_{-p}}\sum_{i\in I_{p}}\bigg{[}||\mu^{t}_{i}(\textbf{h})-\text{Proj}^{t}(\mu^{t}_{p})_{i}(\textbf{h})||^{2}\bigg{]}\right)\leq R^{T}+2.||r||_{\infty}.$
Proof.
By definition of the regret,
/1T$\left(\sum_{t=1}^{T}\varrho_{t}(\textsc{Proj}^{t}(\mu^{t}_{p}))-\varrho_{t}(\mu^{t}_{p})\right)\leq\nicefrac{{1}}{{T}}\left(\max_{\mu\in\tilde{\Lambda}}\sum_{t=1}^{T}\varrho_{t}(\mu_{p})-\varrho_{t}(\mu^{t}_{p})\right)=R^{T}$
Since, $\varrho_{t}(\text{Proj}^{t}(\mu_{p}^{t}))=\mathbb{E}_{\text{Proj}^{t}(\mu_{p}^{t})}[r_{p}(\textbf{h})]$ by property 2, we have,
[…]
which implies the claimed relation. ∎
Appendix E Product games
E.1 General remarks
Here is a slighlty more general presentation of product games. Product games are defined on a product set $\mathbb{H}=\Omega\times\mathbb{U}_{1}\ldots\times\mathbb{U}_{n}$, where the $\mathbb{U}_{i}$ are decision sets. In the present article, those sets are finite. Exogenous randomness is fully encoded in $\Omega$, to which a probability $\mathbb{P}$ is associated. to each decision set $\mathbb{U}_{i}$ is associated a decision maker who’s policy $\mu_{i}$ is a map from $\mathbb{H}$ to $\mathbb{U}_{i}$ satisfying some information constraints, encoded with an information map (like what we do in this article) or a $\sigma$-algebra $\mathcal{I}_{i}$ (the information algebra of decision maker $i$). Without mentioning payoffs, Witsenhausen [45] identifies a condition for this setting to "make sense", i.e. to avoid causal paradox such as $\mu_{1}(h)=1-h_{1}$. He asks for the existence of a solution map $S$ such that for any admissible policy profile $\mu$, $S_{\mu}(\omega)$ is the unique element of $\mathbb{H}$ satisfying $h_{\emptyset}=\omega$ and $\mu_{i}(h)=h_{i}$. This solution map presents several properties related to composition and independence that can be leverage. One can then define the push-forward probability function $\mu\to\mathbb{Q}_{\mu}$ that maps a policy profile to a distribution on $\mathbb{H}$. This generalization of the tree structure allows accounting for situations where the order of play is not clear (typically, when the game can be represented by several trees).
E.2 On lumberjacking
We provide an argument for the claim that modifying information structures can be complex when using trees. Consider any 2-player simultaneous move game, where the players are referred to as player A and player B. To represent this game on a tree, one must decide which player, A or B, plays first. Suppose we decide that A plays before B, but B does not see A’s move. In this setup, it would be possible to extend A’s information by allowing them to see B’s move, but such an extension is not possible with the chosen tree structure. Conversely, using a different tree where B plays before A (and A does not see B’s move) allows for the information extension to make sense. Therefore, the initial tree might not be suitable for information relaxation.
In contrast, in a product game, this can be achieved by manipulating $\mathcal{X}$ without altering $\mathbb{H}$. Note that while we assume a predefined order for simplicity, the assumption of playability used in [15] is much weaker.
Appendix F Details on the experiments
Experiments were performed on a MacBook air laptop with an Apple M1 chip. The code is provided in the supplementary material.
We use $\mathrm{Proj}^{t}=\mathrm{Proj}_{\mu^{t}}$. For each run of the algorithm, the initial state of each learner is sampled with the following code:
weights = 1+np.array([random.random() for _ in range(size)]).astype(np.float64)
weights = weights/ np.sum(weights)
We use the Multiplicative Weights Update (MWU) algorithm as our no-regret learner, with a learning rate set to 1.2 for (2,2) and 0.3 for (3,2). The constants $\lambda$ we use are 0.5 for (2,2) and 3 for (3,2).